OpenStudy (anonymous):

Rewrite in slope-intercept form x=-1/3y-2

OpenStudy (geometry_hater):

slope intecpet y= mx+ b looks like the b is in the right place so how would we get y to the other side?

OpenStudy (anonymous):

y=3x+2

OpenStudy (anonymous):

Divide by y? Can you do that?

OpenStudy (geometry_hater):

Yes you can

OpenStudy (anonymous):

So it's x/y=-1/3-2? That doesn't make sense

OpenStudy (geometry_hater):

Actually you dont even need to move the Y You can just multiply 1/3 (or 3) to the x giving you 3x then add the 2

OpenStudy (anonymous):

ohmygod. thank you

OpenStudy (geometry_hater):

your welcome.